In the laboratory a coffee cup calorimeter, or constant pressure calorimeter, is frequently used to determine the specific
Answer #1

m(water) = 80.36 g

T(water) = 22.73 oC

C(water) = 4.184 J/goC

m(zinc) = 67.53 g

T(zinc) = 98.12 oC

C(zinc) = to be calculated

C3 = 1.61 J/oC

We will be using heat conservation equation

use:

heat lost by zinc = heat gained by water and 3

m(zinc)*C(zinc)*(T(zinc)-T) = m(water)*C(water)*(T-T(water)) + C3*(T-T(water))

67.53*C(zinc)*(98.12-28.0) = 80.36*4.184*(28.0-22.73) + 1.61*(28.0-22.73)

4735.2036*C(zinc) = 1771.9123 + 8.4847

C(zinc)= 0.376 J/goC

Answer: 0.376 J/goC